Complete combustion of 1.0 metric ton of coal (assuming pure carbon) to gaseous carbon dioxide releases 3.3x10^10 J of heat. Convert energy to (a) kilojoules and (b) kilocalorie


1
Expert's answer
2020-10-04T09:16:12-0400

1 kJ = 1000 J

1 cal = 4.184 J

1 BTU = 1055 Joules


1 kcal = 1000 cal


a)

3.3x10^10 J × 1 kJ/1000 J = ? kJ


b)

3.3x10^10 J × 1 cal/4.184 J × 1 kcal/1000 cal = ? kcal


c)

3.3x10^10 J × 1 BTU/1055 J = ? BTU's